# T3-S2S: Training-free Triplet Tuning for Sketch to Scene Generation

### v1

@@ Abstract
Scene generation is crucial to many computer graphics applications. Recent advances in generative AI have streamlined sketch-to-image workflows, easing the workload for artists and designers in creating scene concept art. However, these methods often struggle for complex scenes with multiple detailed objects, sometimes missing small or uncommon instances. In this paper, we propose a Training-free Triplet Tuning for Sketch-to-Scene (T3-S2S) generation after reviewing the entire cross-attention mechanism. This scheme revitalizes the existing ControlNet model, enabling effective handling of multi-instance generations, involving prompt balance, characteristics prominence, and dense tuning. Specifically, this approach enhances keyword representation via the prompt balance module, reducing the risk of missing critical instances. It also includes a characteristics prominence module that highlights TopK indices in each channel, ensuring essential features are better represented based on token sketches. Additionally, it employs dense tuning to refine contour details in the attention map, compensating for instance-related regions. Experiments validate that our triplet tuning approach substantially improves the performance of existing sketch-to-image models. It consistently generates detailed, multi-instance 2D images, closely adhering to the input prompts and enhancing visual quality in complex multi-instance scenes. Code is available at https://github.com/chaos-sun/t3s2s.git.
@@ 1 Introduction
Scene generation plays a significant role in visual content creation across various domains, including video gaming, animation, filmmaking, and virtual/augmented reality. Traditional methods heavily rely on manual efforts, which require designers to transform initial sketches into detailed multi-instance scene concept art through numerous iterations. Recently, technological innovations such as Stable Diffusion [1, 2] equipped with ControlNet [3] and integrated with advanced text-to-image technologies [4], have streamlined this process. These advancements have notably decreased the workload for designers by automating the conversion of simple sketches into complex scenes. While these technologies perform well with common scenes involving typical instances, they struggle with generating complex multi-instance scenes, particularly with unusual and small instances.
Alternatively, multi-instance synthesis involves incorporating layouts of multiple instances as additional input through bounding boxes, and can effectively manage the generation of multiple instances. However, most methods [5, 6, 7, 8, 9, 10] are training-based and require further training when integrated with sketches that contain minimal semantic information, necessitating the collection of numerous scene images. In sectors such as gaming, animation, and film, copyright restrictions significantly hinder scene generation and cannot be disregarded. Conversely, some training-free efforts [11, 12, 13], like Dense Diffusion [4], focus primarily on the impact of the attention map, but they overlook the interaction between the attention map and the value matrices, failing to accurately align with the designer’s sketches.
Our strategy involves maintaining ControlNet’s sketch-following capabilities while exploring the challenges of synthesizing multiple instances. We aim to develop a training-free tuning mechanism that harnesses the inherent creative capabilities of existing models, eliminating the need for extensive data collection or additional training. We conduct a comprehensive analysis of the cross-attention mechanism, identifying two more issues contributing to the performance gaps in generating detailed scenes beyond the attention maps: imbalanced prompt energy and value homogeneity across the cross-attention layers. These two factors often lead to low competitiveness of unusual instances and high coupling among similar instances, resulting in a final image that deviates from the intended instance prompts.
@@
In this paper, we introduce a Training-free Triplet Tuning for Sketch-to-Scene (T3-S2S) generation. Initially, prompt balance improves token representation by adjusting the energy of instance-specific keywords in global text prompts, ensuring that rare instances are adequately represented and remain competitive in the attention mechanism. Subsequently, characteristics prominence distinguishes instance-specific attributes by using a TopK selection strategy from value matrices to amplify feature maps in corresponding channels, highlighting unique instance traits in the multi-channel feature space without extra parameters. Lastly, dense tuning adapted from [4] is applied in the ControlNet branch to refine the contour information of the attention map to compensate for its suboptimal overall strength of instance-related regions. Together, these three tuning strategies form a cohesive triplet strategy that enhances the entire cross-attention mechanism, balancing token competition, enriching the expression of attention maps, and accentuating each instance’s characteristics. Experimental evaluations indicate that our T3-S2S approach boosts the performance of existing text-to-image models, consistently producing detailed, multi-instance scenes that closely align with the input sketches and input prompts.
The key contributions of our work are summarized as follows.
@@
    We investigate the underlying mechanisms of the cross-attention layer and identify the imbalance of prompt energy and homogeneity of value matrices.
@@
    Our T3-S2S model advances a stable diffusion approach by balancing token competition, enriching the expression of attention maps, and accentuating each instance’s characteristics.
@@
    Combined with the triplet tuning, our T3-S2S model enhances the representation of unusual and small instances and realizes high-quality generations of complex multi-instance scenes.
@@
Figure 1: The SDXL-base model [2] and ControlNet model [14] perform well with common instances like humans, but they struggle with complex multi-instance scenes involving small instances and fail to accurately follow users’ prompt.
2 Related Work
Text-to-Image Synthesis. In the rapidly evolving field of text-based image generation, various model architectures and learning paradigms have emerged, as highlighted by several key studies [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Recently, diffusion models  [1, 2, 27] have marked a major breakthrough, significantly improving the fidelity and realism in text-to-image generation, which rely on structured denoising [28] with latent diffusion [1]. Among these, the SDXL model and its variants, which are widely adopted in both academia and industry, are chosen as the baseline for our work.
@@ 2 Related Work
Sketch-to-image Synthesis. While text-to-image models can generate high-fidelity, realistic images, they struggle to accurately convey complex layouts with text prompts alone. In tasks such as scene design for games, animation, film, or virtual reality, hand-drawn sketches with semantic information provide a more effective way to express design ideas. In the field of diffusion-based generation, notable works include ControlNet [3], Make-a-scene [29], and T2I Adapter [30] handle various additional visual conditions, including sketches, while methods like Dense Diffusion [4], SpaText [31] and MultiDiffusion [32] focus specifically on sketch-based inputs. In particular, Dense Diffusion is a training-free approach that adjusts the attention map by amplifying sketch-relevant tokens and downplaying less important ones, allowing the model to better distinguish between instances. ControlNet is a powerful solution for sketch-to-scene generation, recognized for its exceptional ability to accurately follow conditions. However, these models often struggle with complex multi-instance scene generations, particularly when handling unusual or unique instances, and frequently overlook smaller instances. Recently, Sketch2Scene [33] proposed an efficient pipeline for automatically generating interactive 3D game scenes from users’ natural input sketches using the SDXL and ControlNet models. However, the approach is also limited by the diversity and multi-instance representation in the intermediate 2D isometric image generation.
@@
Multi-instance Synthesis. Multi-instance synthesis is closely related to sketch-to-scene generation due to its controllable layout. Training-free modulations  [11, 12, 34, 13] and training-based fine-tuning methods [5, 6, 7, 8, 9, 10] tackle the challenge of diffusion models accurately representing multiple instances with bounding boxes. For example, GLIGEN [6] used bounding box coordinates as grounding tokens, integrating them into a gated self-attention mechanism to improve positioning accuracy, while Detector Diffusion [7] employed a latent object detection model to separate objects, masking conflicting prompts and enhancing relevant ones. Despite existing methods of generating images with correct positions, these box-based approaches struggle with simple sketch inputs and fail to strictly follow the designer’s sketch. Our work leverages ControlNet’s sketch-following capabilities and investigates the challenges of synthesizing multiple instances. We aim to design a training-free tuning mechanism to enhance modeling within cross-attention operations, addressing these challenges effectively.
3 Analyses of Latent Diffusion
3.1 Cross-attention Mechanism
In text-to-image generation tasks, diffusion models aim to transform textual prompts into corresponding images accurately by integrating textual information through cross-attention layers within the UNet model [1, 2]. The mechanism of cross-attention computes attention maps that align intermediate image features with textual embeddings, which can be mathematically represented as:
$\mathbf{F}_{m}=\mathbf{A}_{m}\mathbf{V}_{m}=\text{softmax}\left(\frac{\mathbf{Q}_{m}{\mathbf{K}_{m}}^{T}}{\sqrt{d_{m}}}\right)\mathbf{V}_{m},$ (1)
where $\mathbf{F}_{m}$ is the output of the $m$th cross-attention layer, $\mathbf{A}_{m}$ is the attention map. The query matrices $\mathbf{Q}_{m}\in\mathbb{R}^{{b_{m}}\times d_{m}}$ are derived from the ${m}-1$th intermediate representations within the UNet, where ${b_{m}}$ represents the spatial dimensions (height multiplied by width) and $d_{m}$ is the embedding dimension. The key $\mathbf{K}_{m}\in\mathbb{R}^{n\times d_{m}}$ and value $\mathbf{V}_{m}\in\mathbb{R}^{n\times d_{m}}$ matrices are generated from the encoded text embeddings $\mathbf{S}\in\mathbb{R}^{n\times d}$ from the prompts, where $n$ is the number of tokens and $d$ is the dimension. Based on the mechanism, many explorations [35, 36, 13, 12] and modulations [4, 37, 38, 8] on attention maps tried to figure out how their behaviors at different layers affect the final generations and utilize training-free modulations as well as fine-tuning strategies to improve generation quality. However, despite the significant focus on attention map optimization, there has been relatively little investigation into the entire process of the cross-attention mechanism. Specifically, the role of the $\mathbf{K}_{m}$ matrices in shaping the expression of attention maps, and how $\mathbf{V}_{m}$ contributes to the feature output through the interaction with these maps, remain under-explored.
3.2 Imbalance of Prompt Energy
In extensive practice, two commonly observed phenomena are worth noting: (1) When generating a single instance, the model responds well and rarely misses instances, but in multi-instance generation, some instances are easily lost; (2) In multi-instance generation, if an instance is overlooked, techniques like increasing the prompt weight, such as “(houses:1.5)” in WebUI, can enhance the weight of that prompt after embedding. To quantify this difference, we analyze the text embeddings of multi-instance prompts in Figure 1 (b) and their corresponding single words, using energy (L2 norm) and cosine similarity metrics to measure discrepancies, as shown in Figure 2. For example, in Figure 2(a), the word “houses” exhibits lower energy compared to other words, which may explain why the instance “houses” is easily overlooked in Figure  1 (b). Low energy likely results in lower values in the $\mathbf{K}_{m}$ and $\mathbf{V}_{m}$ matrices during the transformation from text embeddings, leading to diminished attention. In contrast, words encoded separately in Figure 2(b) tend to show higher energy levels, which aligns with the observation that single instances are generally well-represented. Additionally, cosine similarity analysis reveals that embedding full sentences alters the distribution of word importance, reducing the emphasis in some instances. Scaling up the embeddings directly enhances the energy levels, thereby increasing the competitiveness of instances in the generation process by boosting their influence in both the attention map $\mathbf{A}_{m}$ and the value matrix $\mathbf{V}_{m}$. Understanding the imbalance of prompt energy in text embeddings highlights the importance of balancing and scaling energy levels, which offers an interesting perspective to improve multi-instance scene generation.
(a) Multi-instance prompts.
(b) Single-word prompts.
@@ 3.2 Imbalance of Prompt Energy
Figure 2: Comparison of text embeddings between the prompts (“Isometric view of game scene, a plain, walk path, a river, a high mountain, houses.”) and single-word prompts (separate each individual word from the global prompts).
Figure 3: Interaction between attention maps and value matrices with prompts from Figure 2 using dense control [4] in the SDXL-base model and the ControNet model. (a) Sketch-relevant attention map generated by Dense Diffusion. (b) Five-channel value-feature pairs. (c) Two synthetic images were generated with the sketches.
3.3 Homogeneity of Value Matrices
As a key component of cross-attention, the interaction between attention maps and value matrices determines the characteristics of each feature channel related to multiple instances, such as geometry and attributes. However, this process remains poorly understood due to the inherent noise in both attention maps and generated features. Thereby, inspired by Dense Diffusion [4], which enhances sketch-relevant values of the attention map $\mathbf{A}_{m}\in\mathbb{R}^{{b_{m}}\times n}$ depicted in Figure 3 (a), this strategy effectively highlights different instances for each token with a defined level of emphasis. Then, we visualize five-channel value-feature pairs from the $\{\mathbf{v}^{j}_{m}\in\mathbb{R}^{n}\}^{d_{m}}_{j=1}$ (denoted as $\mathbf{V}_{m}$) and corresponding feature map $\{\mathbf{f}^{j}_{m}\in\mathbb{R}^{{b_{m}}}\}^{d_{m}}_{j=1}$ (denoted as $\mathbf{F}_{m}$) in Figure 3 (b).
@@ 3.3 Homogeneity of Value Matrices
Figure 4: Generations by amplifying the TopK values of the value matrices based on the pipeline in Figure 3.
From Figure 3, we can observe: (1) Extremums matter: Tokens with extreme values far from zero generate stronger instance characteristics when interacting with attention maps. (2) Small areas overlooked: Instances with small areas, such as “Path” and “Houses”, are easily neglected in the final image, despite having strong responses in feature maps. This can resemble homogeneity of values, where numerical differences between tokens in the value matrix are minimal, and the model struggles to distinguish between instances, leading to instance coupling and the failure to generate certain instances in the final image. This highlights the need for significant numerical disparities among tokens to ensure instance representation.
@@
To assess this potential, we amplify the TopK values in each channel of the value matrices two-fold, as shown in Figure 4. As K increases, particularly at K=2, the model initially generates all instances successfully. However, this also introduces excessive noise, cluttering images with unnecessary details, as seen in the depiction of houses. This experiment suggests that increasing the TopK values enhances token competitiveness and reduces value homogeneity. However, it also highlights a trade-off between instance completeness and visual clarity, underscoring the need for a balanced approach to value amplification in dense diffusion.
4 PROPOSED APPROACH
In this section, we present an overview outlining the comprehensive mechanism of our approach. This is followed by a detailed examination of the prompt balance and the characteristics prominence.
4.1 Overview
To address the challenges of the cross-attention mechanism highlighted in Section 3, we introduce the training-free triplet tuning strategy, which builds on the strengths of the pre-trained SDXL and ControlNet models, incorporating textual prompts $\mathbf{c}_{g}=\{c^{i}\}^{l}_{i=0}$ ($l$ is the number of words) and corresponding sketch images $\mathbf{C}_{s}\in\mathbb{R}^{h\times w}$, as detailed in Figure 5.
The proposed training-free triplet tuning can be divided into the following three modules.

### v2

# T3-S2S: Training-free Triplet Tuning for Sketch to Scene Synthesis in Controllable Concept Art Generation

@@ Abstract
2D concept art generation for 3D scenes is a crucial yet challenging task in computer graphics, as creating natural intuitive environments still demands extensive manual effort in concept design. While generative AI has simplified 2D concept design via text-to-image synthesis, it struggles with complex multi-instance scenes and offers limited support for structured terrain layout. In this paper, we propose a Training-free Triplet Tuning for Sketch-to-Scene (T3-S2S) generation after reviewing the entire cross-attention mechanism. This scheme revitalizes the ControlNet model for detailed multi-instance generation via three key modules: Prompt Balance ensures keyword representation and minimizes the risk of missing critical instances; Characteristic Priority emphasizes sketch-based features by highlighting TopK indices in feature channels; and Dense Tuning refines contour details within instance-related regions of the attention map. Leveraging the controllability of T3-S2S, we also introduce a feature-sharing strategy with dual prompt sets to generate layer-aware isometric and terrain-view representations for the terrain layout. Experiments show that our sketch-to-scene workflow consistently produces multi-instance 2D scenes with details aligned with input prompts. Code is available at the official repository https://github.com/Tencent/Triplet_Tuning, with a maintained mirror for future updates at https://github.com/EngineeringAI-LAB/triplet_tuning.
@@ 1 Introduction
Scene generation plays a significant role in visual content creation across various domains, including video gaming, animation, filmmaking, and virtual/augmented reality. Traditional methods heavily rely on manual efforts, which require designers to transform initial sketches into detailed multi-instance scene concept art through numerous iterations. Recently, technological innovations such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) equipped with ControlNet (Zhang et al., 2023b) and integrated with advanced text-to-image technologies (Kim et al., 2023), have streamlined this process. These advancements have notably decreased the workload for designers by automating the conversion of simple sketches into complex scenes. While these technologies perform well with common scenes involving typical instances, they struggle with generating complex multi-instance scenes and often miss fine details. For the terrain layout, designers still manually interpret concept art and build terrain by dragging grids with preset tools.
Alternatively, sketch-based synthesis can adopt multi-instance strategies by incorporating instance layouts via bounding boxes to guide the generation of multiple elements. While effective, most existing methods (Yang et al., 2023; Li et al., 2023; Liu et al., 2023; Sun et al., 2024; Wang et al., 2024b; Zhou et al., 2024) are training-based and require adaptation for sketches, relying on large datasets often restricted by copyright in gaming, animation, and film. In contrast, training-free approaches (Xie et al., 2023; Chen et al., 2024; Feng et al., 2022; Kim et al., 2023) leverage attention maps to exploit inherent model capabilities, offering flexible adaptability and low cost for new tasks. Building on this, we retain ControlNet’s sketch-following ability with a training-free tuning mechanism requiring no additional data. Through a theoretical cross-attention analysis, we find that imbalanced prompt energy and value non-prominence undermine the competitiveness of instances and increase coupling among similar ones, leading to deviation from intended prompts.
@@
In this paper, we introduce a Training-free Triplet Tuning for Sketch-to-Scene (T3-S2S) generation via three modules. Prompt balance adjusts instance-specific keyword energy to ensure all instances remain competitive. Characteristics priority amplifies instance-specific traits using a TopK selection from value matrices to enhance channel prominence in the feature map. Dense tuning adapted from (Kim et al., 2023) strengthens instance-related contour details in the attention map of the ControlNet branch. Together, these three form a unified triplet strategy that generates detailed, multi-instance 2D scenes that closely align with input prompts and sketches. Based on the controllability, we propose a twin structural concept framework that uses a dual-branch prompt synthesis with a mask-guided feature-sharing mechanism to generate layer-aware isometric and terrain-view representations for reconstructing the terrain layout. Experimental evaluations indicate that our T3-S2S approach boosts the performance of existing text-to-image models, consistently producing detailed, multi-instance scenes that closely align with the input sketches and input prompts.
The key contributions of our work are summarized as follows:
@@
    We theoretically investigate the underlying cross-attention mechanisms and identify the imbalance of prompt energy and value non-prominence, leading to deviation from intended prompts.
@@
    The triplet tuning advances controllable concept art generation by balancing token competition, enriching attention expression, and accentuating each instance’s characteristics.
@@
    Our T3-S2S  workflow consistently generates detailed multi-instance 2D images and terrain layout aligned with input prompts.
@@
Figure 1: The SDXL-base model (Podell et al., 2023) and ControlNet model (Xinsir, 2023) perform well with common instances like humans, but they struggle with complex multi-instance scenes involving small instances and fail to accurately follow users’ prompt.
2 Related Works
Diffusion Techniques. Recently, diffusion models  (Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023; Saharia et al., 2022) have marked a major breakthrough, improving the fidelity and realism in text-to-image generation, which shows infinite potential for concept art generation. The emergence of diffusion models has also advanced the development of 3D content generation tools. However, creating high-fidelity 3D scenes from images remains a complex task due to the diversity and intricacy of object shapes and appearances.
@@ 2 Related Works
Sketch-to-image Synthesis. While text-to-image models can generate high-fidelity, realistic images, they struggle to accurately convey complex layouts with text prompts alone. In the field of diffusion-based generation, notable works include ControlNet (Zhang et al., 2023b), Make-a-scene (Gafni et al., 2022), and T2I Adapter (Mou et al., 2023) handle various additional visual conditions, including sketches, while methods like Dense Diffusion (Kim et al., 2023), SpaText (Avrahami et al., 2023) and MultiDiffusion (Bar-Tal et al., 2023) focus specifically on sketch-based inputs. In particular, Dense Diffusion is a training-free approach that adjusts the attention map by amplifying sketch-relevant tokens and downplaying less important ones, allowing the model to better distinguish between instances. ControlNet is a powerful solution for sketch-to-scene generation, recognized for its exceptional ability to accurately follow conditions. However, these models often struggle with complex multi-instance scene generations, particularly when handling unusual or unique instances, and frequently overlook smaller instances. Recently, Xu et al. (2024) proposed an efficient pipeline for automatically generating interactive 3D game scenes from users’ natural input sketches using the SDXL and ControlNet models. However, the approach is also limited by the diversity and multi-instance representation in the intermediate 2D isometric image generation.
Figure 2: The SDXL-base (Podell et al., 2023) and ControlNet models (Xinsir, 2023) struggle with complex multi-instance scenes based on sketch images and text prompt, even with improved dense tuning (Kim et al., 2023).
@@
Multi-instance Synthesis. Multi-instance synthesis is closely related to sketch-to-scene generation due to its controllable layout. Training-free modulations  (Xie et al., 2023; Chen et al., 2024; Lian et al., 2023; Feng et al., 2022) and training-based fine-tuning methods (Yang et al., 2023; Li et al., 2023; Liu et al., 2023; Sun et al., 2024; Wang et al., 2024b; Zhou et al., 2024) tackle the challenge of diffusion models accurately representing multiple instances with bounding boxes. For example, Li et al. (2023) (GLIGEN) used bounding box coordinates as grounding tokens, integrating them into a gated self-attention mechanism to improve positioning accuracy, while Liu et al. (2023) employed a latent object detection model to separate objects, masking conflicting prompts and enhancing relevant ones. Despite existing methods of generating images with correct positions, these box-based approaches struggle with simple sketch inputs and fail to strictly follow the designer’s sketch. Our work leverages ControlNet’s sketch-following capabilities and investigates the challenges of synthesizing multiple instances. We aim to design a training-free tuning mechanism to enhance modeling within cross-attention, addressing these challenges effectively.
3 Controllability Analysis of Cross-attention Mechanism
Problem Statement. ControlNet (Zhang et al., 2023a) is adopted alongside SDXL (Podell et al., 2023) as our baseline for concept art generation due to its controllability in sketch-to-image generation. Given a textual prompt $\mathbf{c}_{g}=\{c^{i}\}_{i=0}^{l}$ ($l$ is words number) and a sketch image $\mathbf{C}_{s}\in\mathbb{R}^{h\times w}$, the system generates images via cross-attention mechanism in the UNet (Rombach et al., 2022), aligning spatial features with text embeddings $\mathbf{S}\in\mathbb{R}^{n\times d}$ (encoded from $\mathbf{c}_{g}$):
$\mathbf{F}_{m}=\mathbf{A}_{m}\mathbf{V}_{m}=\text{softmax}\left({\mathbf{Q}_{m}\mathbf{K}_{m}^{\top}}/{\sqrt{d_{m}}}\right)\mathbf{V}_{m},$
where $m$ is layer number, $n$ the number of text tokens, and $d$ the token embedding dimension. $\mathbf{Q}_{m}\in\mathbb{R}^{b_{m}\times d_{m}}$ comes from fore-features, and $\mathbf{K}_{m},\mathbf{V}_{m}\in\mathbb{R}^{n\times d_{m}}$ are derived from text embeddings $\mathbf{S}$ by linear projections, where $b_{m}$ is the flattened spatial dimension, and $d_{m}$ is the embedding dimension. As shown in Figure 2, the system struggles with complex multi-instance natural scenes guided by sketch images and text prompts, often failing to render all intended objects.
Dense Diffusion (Kim et al., 2023) can effectively highlight different instances’ attention values based on the sketches in the attention map $\mathbf{A}_{m}\in\mathbb{R}^{b_{m}\times n}$, but directly applying its strategy to ControlNet leads to performance degradation. To adapt it, we make an improved Dense Tuning (DT) (Details in Appendix B): 1) Only add the dense tuning at the “down_block_2" and “mid_block_0" of ControlNet; 2) Only use the expand function, not use the suppress function in the self-attention. Despite this, the image quality improves, but instances with small areas, such as “path" and “houses", are easily neglected in the final image, despite having strong responses in feature maps. This highlights that beyond tuning attention maps alone, there remains room to explore controllability within the broader cross-attention process. In particular, the roles of $\mathbf{K}_{m}$ in shaping attention distributions and $\mathbf{V}_{m}$ in contributing to the final feature outputs have received limited attention.
(a) Single-word prompts
(b) Multi-word prompts
@@ 3 Controllability Analysis of Cross-attention Mechanism
Figure 3: Embedding energy comparison between a global prompt (“Isometric view of game scene, a plain, walk path, a river, a high mountain, houses.”) and single-word prompts (each keyword separated and embedded individually, higher than that in the group). The energy imbalance can lead to attention competition, and low-energy small instances (“path" and “houses") are easily forgotten. (c) Cosine similarity between embeddings of (a) and (b)
Imbalance of Prompt Energy. In practice, increasing prompt weights (e.g., “(houses:1.5)” in WebUI) can enhance the visibility of specific instances in multi-instance generation. To understand the underlying mechanism, we refer to Henry et al. (2020) to decompose each token embedding from $\mathbf{S}=[\mathbf{s}^{1},\dots,\mathbf{s}^{n}]\in\mathbb{R}^{n\times d}$ as:
$\mathbf{s}^{i}=E^{i}\cdot\hat{\mathbf{s}}^{i},\quad\text{where }E^{i}=\|\mathbf{s}^{i}\|,\ \|\hat{\mathbf{s}}^{i}\|=1,$
where, $E_{i}$ represents the prompt energy (L2 norm), and $\hat{\mathbf{s}}^{i}$ denotes the normalized embedding. Since both keys and values in cross-attention are linear projections of $\mathbf{s}^{i}$, their magnitudes are bounded:
$\|\mathbf{k}^{i}\|=\|\mathbf{s}^{i}\mathbf{W}_{K}\|\leq E^{i}\cdot\|\mathbf{W}_{K}\|,\quad\|\mathbf{v}^{i}\|=\|\mathbf{s}^{i}\mathbf{W}_{V}\|\leq E^{i}\cdot\|\mathbf{W}_{V}\|.$
Thus, tokens with lower energy yield weaker keys $\mathbf{k}^{i}$ and values $\mathbf{v}^{i}$, diminishing their impact on attention and feature aggregation.
Considering the prompts in Fig. 2, we analyze the embedding energy within a global prompt and each keyword embedding separately, as shown in Fig. 3. Global encoding obviously lowers the energy of the individual tokens, especially the instances like “path” and “houses”, aligns with the observed instance omission in Fig. 2. The final attention value $a$ of the valid token can be represented as $a\approx E\bar{a}$, where $\bar{a}$ denotes the mean attention weight, empirically observed as approximately $0.43$ across $100$ prompt sets. After applying exponentiation, the ratio between two attention values becomes:
$Ratio={e^{E^{2}\bar{a}}}/{e^{E^{1}\bar{a}}}=e^{(E^{2}-E^{1})\bar{a}}.$
Due to exponential weighting, energy differences of $\{1,3,5\}$ result in an attention $Ratio$ disparity of $\{1.54,3.63,8.58\}$ before the softmax average, and higher-energy tokens easily win the attention competition. Even if a low-energy token receives high attention via tuning, its contribution remains limited due to the low magnitude of its value vector $\mathbf{v}_{i}$ which weakens the aggregated feature in the product $\mathbf{A}_{m}\mathbf{V}_{m}$. This energy imbalance, which allows dominant tokens to overshadow weaker ones and increases the risk of instance omission in multi-instance synthesis, underscores the importance of balancing and scaling prompt energy as an interesting perspective to improve multi-instance scene generation.
Figure 4: Interaction between attention maps and value matrices with prompts from Fig. 2 using dense tuning. (a) Attention maps highlight strong sketch relevance. (b/c) Five-channel value-feature pairs reveal the importance of extrema. Despite feature enhancement improving instance chances, forgetting still occurs as extrema are not prominent. Statistics are shown in Fig. 10 and Appendix C.
@@
Figure 5: Generations by amplifying the TopK extrema twice in the value matrices based on the pipeline in Fig. 5, where most instances appear, but uniform amplification also introduces noise.
Non-prominence of Value Matrices. As a core part of cross-attention, the interaction between attention maps and value matrices shapes each channel’s characteristics to instance-level patterns such as geometry and attributes. To maximize expressiveness, each channel should focus on distinct semantics, similar to SENet’s calibration (Hu et al., 2018). The aggregated feature at channel $j$ is:
$\mathbf{f}^{j}_{m}=\mathbf{A_{m}}\cdot\mathbf{v}^{j}_{m}=\{[a^{z,1}_{m},\ a^{z,i}_{m},\ \dots,\ a^{z,n}_{m}]\cdot[v^{1,j}_{m},\ v^{i,j}_{m},\ \dots,\ v^{n,j}_{m}]\}^{b_{m}}_{z=1}.$
As shown in Fig. 5(a), dense tuning sharpens attention, allowing each spatial position $z$ to attend to a single instance via high $a^{z,i}_{m}$. However, The non-prominence between value values can cause responses to multiple instances, leading to entangled features. Specifically, the extreme values determine which instances are activated spatially, consistent with the observations in Fig. 5(b) and (c).
@@
To probe this, we double the TopK values in each channel of the value matrices, as shown in Figure 5. As $K$ increases, the model initially synthetizes all instances (e.g., at $K=2$), but it also introduces excessive noise (e.g., over-detailed houses). This demonstrates that stronger value prominence improves token competitiveness, but requires a trade-off between instance completeness and visual clarity via controlled value amplification.
4 Diffusion for Controllable Concept Art Generation
To address the cross-attention challenges discussed in Section 3, we propose a training-free triplet tuning strategy for controllable concept art generation, as illustrated in Fig. 6. This strategy consists of three modules: Prompt Balance (PB), Characteristics Priority (CP), and Dense Tuning (DT):